# Note about non-compact embeddings between Besov spaces

### v1

@@ 3. Embeddings Between Sequence Spaces
  • (Propositions 3.7 and 3.8) Assume either $p_{1}<p_{0}<\infty$ or $p_{0}=p_{1}=\infty$. Then,
@@
[…]
@@
By assumption, for every $n$-dimensional subspace $V\subset\ell^{p}(X_{j})_{j=0}^{\infty}$ there is a unit vector $x\in V$ such that $\|\widehat{T}_{p,\infty}x\|_{\ell^{\infty}(Y_{j})_{j}}\leq b_{n}(\widehat{T}_{p,\infty})\|x\|_{\ell^{p}(X_{j})_{j}}$. Therefore $\|\widehat{T}_{p,q}x\|_{\ell^{q}(Y_{j})_{j}}\leq\|\widehat{T}_{p,\infty}x\|_{\ell^{\infty}(Y_{j})_{j}}^{1-p/q}\leq b_{n}(\widehat{T}_{p,\infty})^{1-p/q}$. Since $\widehat{T}_{p,\infty}$ is finitely strictly singular by the above, the right hand side tends to 00, which concludes the proof. ∎
@@
The key to prove Theorem 1.2 ii is the following
@@
This can be a direct corollary if we have the quasi-Banach version of the diagonal theorem. For completeness we provide a proof following the similar argument to Corollary 3.4.
@@
The Hölder’s inequality yields $\|x\|_{\ell^{q/\theta}(\ell^{p/\theta})}\leq\|x\|_{\ell^{q}(\ell^{p})}^{\theta}\|x\|_{\ell^{\infty}}^{1-\theta}$ for all $0<p,q\leq\infty$ and $10<\theta<$ 1. Therefore,
@@
Let $0<q_{0}<q_{1}\leq\infty$ and $0<p\leq\infty$. Let $(m_{j})_{j=0}^{\infty}\subset\mathbb{Z}_{+}$ be an arbitrary sequence.
@@
$\ell^{q_{0}}(\ell^{p}\{1,\dots,m_{j}\})_{j=0}^{\infty}\hookrightarrow\ell^{q_{1}}(\ell^{p}\{1,\dots,m_{j}\})_{j=0}^{\infty}.$
When $\sup m_{j}=\infty$ and $q_{1}<\infty$ the map is not finitely strictly singular. For the special case $m_{j}=2^{j}$, see Proposition 3.8 below.
@@
Let $\varepsilon>0$ and $V\subset\ell^{q_{0}}(\ell^{p}\{1,\dots,m_{j}\})_{j=0}^{\infty}$ be an infinite dimensional subspace. We need to find a non-zero $x=(x_{j})_{j=0}^{\infty}\in V$ such that $\|x\|_{\ell^{q_{1}}(\ell^{p})}<\varepsilon\|x\|_{\ell^{q_{0}}(\ell^{p})}$.
@@
Since $q_{0}<\infty$ and $\{1,\dots,m_{j}\}$ are finite sets, we can find out a sequence $(j_{n})_{n=1}^{N}\subset\mathbb{N}_{0}$ and a $(x^{(n)})_{n=1}^{N}\subset V$ with $\|x^{(n)}\|_{\ell^{q_{0}}(\ell^{p})}=1$ for each $n\geq 1$, such that
@@
  • $\sum_{j=j_{n}+1}^{\infty}\|x^{(n)}_{j}\|_{\ell^{p}\{1,\dots,m_{j}\}}^{q_{0}}\leq\delta$, or equivalently, $\sum_{j=j_{n-1}+1}^{j_{n}}\|x^{(n)}_{j}\|_{\ell^{p}\{1,\dots,m_{j}\}}^{q_{0}}\geq 1-\delta$.
The construction is done by induction. We start with a unit vector $x^{(1)}\in V$, since $q_{0}<\infty$ there is a $j_{1}\geq 1$ such that $\sum_{j=j_{1}+1}^{\infty}\|x^{(1)}_{j}\|_{\ell^{p}\{1,\dots,m_{j}\}}^{q_{0}}\leq\delta$.
Suppose for $n\geq 2$ and we have obtained $x^{(1)},\dots,x^{(n-1)}$ and $j_{1}<\dots<j_{n-1}$. Consider the projection $\mathcal{P}_{n-1}:\ell^{q_{0}}(\mathbb{N}_{0};\ell^{p})\twoheadrightarrow\ell^{q_{0}}(\{0,\dots,j_{n-1}\};\ell^{p})$, we have $\operatorname{ker}(\mathcal{P}_{n-1}|_{V})\neq\{0\}$ since $\dim\mathcal{P}_{n-1}(V)=\sum_{j=0}^{j_{n-1}}m_{j}<\infty$ but $\dim V=\infty$. Therefore, there must be a unit vector $x^{(n)}\in V$ such that $\mathcal{P}_{n-1}x^{(n)}=0$, i.e. $x^{(n)}_{j}\equiv 0$ for every $0\leq j\leq j_{n-1}$. Take a $j_{n}>j_{n-1}$ be such that $\sum_{j=j_{n}+1}^{\infty}\|x^{(n)}_{j}\|_{\ell^{p}\{1,\dots,m_{j}\}}^{q_{0}}\leq\delta$ we close the induction.
@@
$\|x\|_{\ell^{q_{0}}(\ell^{p})}\geq\|x\|_{\ell^{q_{0}}(\{1,\dots,j_{1}\};\ell^{p})}=\|x^{(1)}\|_{\ell^{q_{0}}(\{1,\dots,j_{1}\};\ell^{p})}\geq(1-\delta)^{1/q_{0}}>0.$
@@
$\Big{\|}\sum_{k=1}^{N}x^{(k)}\Big{\|}_{\ell^{q_{0}}(\{j_{n-1}+1,\dots,j_{n}\};\ell^{p})}^{\min(1,q_{0})}$ $=\Big{\|}\sum_{k=1}^{n}x^{(k)}\Big{\|}_{\ell^{q_{0}}(\{j_{n-1}+1,\dots,j_{n}\};\ell^{p})}^{\min(1,q_{0})}$
$\geq\|x^{(n)}\Big{\|}_{\ell^{q_{0}}(\{j_{n-1}+1,\dots,j_{n}\};\ell^{p})}^{\min(1,q_{0})}-\Big{\|}\sum_{k=1}^{n-1}x^{(k)}\Big{\|}_{\ell^{q_{0}}(\{j_{n-1}+1,\dots,j_{n}\};\ell^{p})}^{\min(1,q_{0})}$
$\geq\|x^{(n)}\Big{\|}_{\ell^{q_{0}}(\{j_{n-1}+1,\dots,j_{n}\};\ell^{p})}^{\min(1,q_{0})}-\sum_{k=1}^{n-1}\|x^{(k)}\|_{\ell^{q_{0}}(\{j_{k}+1,j_{k}+2,\dots\};\ell^{p})}^{\min(1,q_{0})}$
@@
$\Big{\|}\sum_{k=1}^{N}x^{(k)}\Big{\|}_{\ell^{q_{0}}(\ell^{p})}^{q_{0}}\geq\sum_{n=1}^{N}\Big{\|}\sum_{k=1}^{N}x^{(k)}\Big{\|}_{\ell^{q_{0}}(\{j_{n-1}+1,\dots,j_{n}\};\ell^{p})}^{q_{0}}\geq N(1-(N+1)\delta^{\min(1,q_{0}^{-1})})^{\frac{q_{0}}{\min(1,q_{0})}}.$
On the other hand, using the property $\|x\|_{\ell^{q_{1}}(S;\ell^{p})}\leq\|x\|_{\ell^{q_{0}}(S;\ell^{p})}$ for every $x\in\ell^{q_{0}}(\ell^{p})$ and set $S$, we have, for each $1\leq n\leq N$,
$\Big{\|}\sum_{k=1}^{N}x^{(k)}\Big{\|}_{\ell^{q_{1}}(\{j_{n-1}+1,\dots,j_{n}\};\ell^{p})}^{\min(1,q_{1})}=$ $\Big{\|}\sum_{k=1}^{n}x^{(k)}\Big{\|}_{\ell^{q_{1}}(\{j_{n-1}+1,\dots,j_{n}\};\ell^{p})}^{\min(1,q_{1})}$
$\leq\|x^{(n)}\Big{\|}_{\ell^{q_{1}}(\{j_{n-1}+1,\dots,j_{n}\};\ell^{p})}^{\min(1,q_{1})}$ $+\Big{\|}\sum_{k=1}^{n-1}x^{(k)}\Big{\|}_{\ell^{q_{1}}(\{j_{n-1}+1,\dots,j_{n}\};\ell^{p})}^{\min(1,q_{1})}$
$\leq\|x^{(n)}\Big{\|}_{\ell^{q_{0}}(\{j_{n-1}+1,\dots,j_{n}\};\ell^{p})}^{\min(1,q_{1})}$ $+\sum_{k=1}^{n-1}\|x^{(k)}\|_{\ell^{q_{1}}(\{j_{k}+1,j_{k}+2,\dots\};\ell^{p})}^{\min(1,q_{1})}\leq 1+N\delta^{\frac{\min(1,q_{1})}{q_{0}}}.$
@@
$\Big{\|}\sum_{k=1}^{N}x^{(k)}\Big{\|}_{\ell^{q_{1}}(\{j_{N}+1,j_{N}+2,\dots\};\ell^{p})}^{\min(1,q_{1})}\leq\sum_{k=1}^{N}\|x^{(k)}\|_{\ell^{q_{1}}(\{j_{N}+1,j_{N}+2,\dots\};\ell^{p})}^{\min(1,q_{1})}\leq\delta^{\frac{\min(1,q_{1})}{q_{0}}}N.$
@@
$\Big{\|}\sum_{k=1}^{N}x^{(k)}\Big{\|}_{\ell^{q_{1}}(\ell^{p})}\leq\bigg{(}\sum_{n=1}^{N+1}\Big{\|}\sum_{k=1}^{N}x^{(k)}\Big{\|}_{\ell^{q_{1}}(\{j_{n-1}+1,\dots,j_{n}\};\ell^{p})}^{q_{1}}\bigg{)}^{1/q_{1}}\leq N^{1/q_{1}}(1+(N+1)\delta^{\frac{\min(1,q_{1})}{q_{0}}})^{\frac{1}{\min(1,q_{1})}}.$
@@
$\frac{\|x^{(1)}+\dots+x^{(N)}\|_{\ell^{q_{1}}(\ell^{p})}}{\|x^{(1)}+\dots+x^{(N)}\|_{\ell^{q_{0}}(\ell^{p})}}\leq N^{\frac{1}{q_{1}}-\frac{1}{q_{0}}}\frac{(1+(N+1)\delta^{\frac{\min(1,q_{1})}{q_{0}}})^{\frac{q_{1}}{\min(1,q_{1})}}}{(1-(N+1)\delta^{\min(1,\frac{1}{q_{0}})})^{\frac{q_{0}}{\min(1,q_{0})}}}<\frac{\varepsilon}{2}\cdot 2=\varepsilon.$
@@
When $p_{0}=p_{1}$, we can take the $2^{n}$ dimensional subspace $V_{n}:=2^{-n/{p_{0}}}\cdot\ell^{p_{0}}\{1,\dots,2^{n}\}\hookrightarrow\ell^{q_{0}}(2^{-j/{p_{0}}}\cdot\ell^{p_{0}}\{1,\dots,2^{j}\})_{j=0}^{\infty}$, where $n\geq 1$. The embedding (3.3) restricted to $V_{n}$ is the identity map, hence the Bernstein numbers are all bounded from below by $1$, which does not tend to 00.
@@
Since $(S_{n},\mu_{n})$ and ($(\{1,\dots,2^{n}\},2^{-n}\cdot\#)$ are both probability spaces with uniform distributions on their sample sets, we see that they are isomorphic by taking any bijection $\Gamma_{n}:\{1,\dots,2^{n}\}\to S_{n}$ between sets. Thus we have elements $r_{n,j}\circ\Gamma_{n}\in\ell^{p_{1}}\{1,\dots,2^{n}\}$.
@@
    Suppose $q_{0}<q_{1}$, and either $p_{1}<p_{0}<\infty$ or $p_{0}=p_{1}=\infty$, then the embedding is strictly singular but not finitely strictly singular.
@@
The map $J_{s}(x_{j})_{j=A}^{\infty}:=(2^{js}x_{j})_{j=A}^{\infty}$ defines an isomorphism $J_{s}:\ell^{q}(2^{j(s-\frac{n}{p})}\cdot\ell^{p}\{1,\dots,2^{nj}\})_{j=A}^{\infty}\to\ell^{q}(2^{-j\frac{n}{p}}\cdot\ell^{p}\{1,\dots,2^{nj}\})_{j=A}^{\infty}$ for every $0<p,q\leq\infty$. Therefore it suffices to consider the case $s=0$ for both results.
i: For $j\geq 0$ and $f:\{1,\dots,2^{nj}\}\to\mathbb{R}$ we define $\Gamma_{j}f:[0,1]\to\mathbb{R}$ by $\Gamma_{j}f(x):=f(\lfloor 2^{nj}x\rfloor)$. Clearly $\Gamma_{j}:2^{-jn/p}\cdot\ell^{p}\{1,\dots,2^{nj}\}\to L^{p}[0,1]$ are isometric embeddings for all $0<p\leq\infty$ and $j\geq 0$. Therefore $(\Gamma_{j})_{j=A}^{\infty}:\ell^{q}(2^{-jn/p}\cdot\ell^{p}\{1,\dots,2^{nj}\})_{j=A}^{\infty}\to\ell^{q}(L^{p}[0,1])$ is an isometric embedding for all $0<p,q\leq\infty$ and $A\geq 1$.
@@
ii: The non finitely strictly singular part follows from Proposition 3.8 while the strictly singular part follows from Proposition 3.7.
@@
[…]
@@
Since $\iota_{0},\iota_{1}$ are both topological embedding, there is a $C>0$ such that $C^{-1}\|x\|_{X_{k}}\leq\|\iota_{k}x\|_{Y_{k}}\leq C\|x\|_{X_{k}}$ for all $x\in X_{k}$, $k=0,1$. Both results on the finitely strictly singularity follow from the fact that $b_{n}(T_{X})\leq C^{2}b_{n}(T_{Y})$. The strictly singularity follows from the same argument and we leave the details to the reader. ∎
@@
ii: When $s_{0}-s_{1}=\frac{n}{p_{0}}-\frac{n}{p_{1}}$ and $q_{0}=q_{1}$, we are going to find an infinite dimensional subspace $X$ such that the embedding restricted to $X$ has bounded inverse.
Let $A\geq 1$ be as in Proposition 2.15. Then, there exists $m\in Q_{A}$ such that $\operatorname{supp}\psi_{jm}^{g}\subset\Omega$ for all $j\geq A$ and $g\in G_{j}(=\{0,1\}^{n}\backslash\{0\}^{n})$. We fix a $g\in\{0,1\}^{n}\backslash\{0\}^{n}$ and consider $X:=\operatorname{span}\{\psi_{jm}^{g}:j\geq A\}$. We see that $X$ is infinite dimensional and is a subspace of both $\mathscr{B}_{p_{0}q_{0}}^{s_{0}}(\Omega)$ and $\mathscr{B}_{p_{0}q_{0}}^{s_{0}}(\mathbb{R}^{n})$. By Proposition 2.12 iv we have, for every $(a_{j})_{j=A}^{\infty}\subset\mathbb{R}$,
@@
Proposition 4.4.
@@
Still we take an $\varepsilon>0$ be such that $p_{0},q_{0},p_{1},q_{1}>\varepsilon$ and $|s_{0}|,|s_{1}|<\varepsilon^{-1}$. Let $A\geq 1$ and $R_{j},S_{j}\subset Q_{j}\times G_{j}$ for $j\geq A$ be from Proposition 2.15 where we choose fixed subsets $V\Subset\Omega\Subset U$. Recall that #$\#R_{j}=2^{(j-A)n}$ for $j\geq A$, #$\#S_{j}=2^{(j+A)n}$ for $j\geq 0$, and for $k=0,1$ both arrows in $\ell^{q_{k}}(2^{j(s-\frac{n}{p_{k}}+\frac{n}{2})}\cdot\ell^{p_{k}}(R_{j}))_{j=A}^{\infty}\xrightarrow{\Lambda^{-1}}\mathscr{B}_{p_{k}q_{k}}^{s}(\Omega)\xrightarrow{\Lambda\circ E}\ell^{q_{k}}(2^{j(s_{k}-\frac{n}{p_{k}}+\frac{n}{2})}\cdot\ell^{p_{k}}(S_{j}))_{j=0}^{\infty}$ are bounded.
@@
iii: By Corollary 2.17 ii the map is noncompact. For finitely strictly singularity, when $s_{0}-s_{1}=\frac{n}{p_{0}}-\frac{n}{p_{1}}$ the result follows from Proposition 4.2 i; when $s_{0}-s_{1}=0$ the result follows from Proposition 4.4 i.
@@
v: When $s_{0}-s_{1}=\frac{n}{p_{0}}-\frac{n}{p_{1}}$ the result follows from Proposition 4.2 ii; when $s_{0}-s_{1}=0$ the result follows from Proposition 4.4 iii. ∎

### v2

@@ 3. Embeddings Between Sequence Spaces
  • (Propositions 3.7 and 3.8) Assume either $p_{1}\leq p_{0}<\infty$ or $p_{0}=p_{1}=\infty$. Then,
@@
These results are inspired from Plichko [Pli04] and Lefèvre, Rodríguez-Piazza [LRP14]. Although some of the techniques might not be new, to the best of the authors’ knowledge, the proof is not found in the standard literature. Thus we attach a full proof in detail.
@@
Let $e_{j}=(0,\dots,0,\underset{j}{1},0,\dots)\in\ell^{\infty}$ be the standard elements. By assumption, there exist $c>0$ and $x_{j}\in X_{j}$ such that for all $j\geq 0$, $\|x_{j}\|_{X_{j}}=1$ and $\|T_{j}x_{j}\|_{Y_{j}}\geq c$.
@@
By assumption, for every $n$-dimensional subspace $V\subset\ell^{p}(X_{j})_{j=0}^{\infty}$ there is a unit vector $x\in V$ such that $\|\widehat{T}_{p,\infty}x\|_{\ell^{\infty}(Y_{j})_{j}}\leq b_{n}(\widehat{T}_{p,\infty})\|x\|_{\ell^{p}(X_{j})_{j}}$. Therefore $\|\widehat{T}_{p,q}x\|_{\ell^{q}(Y_{j})_{j}}\leq\|\widehat{T}_{p,\infty}x\|_{\ell^{\infty}(Y_{j})_{j}}^{1-p/q}\leq b_{n}(\widehat{T}_{p,\infty})^{1-p/q}$. Since $\widehat{T}_{p,\infty}$ is finitely strictly singular by the above, the right hand side tends to 0, which concludes the proof. ∎
@@
The proof of Theorem 1.2 ii is reduced to the following statement on the sequence spaces.
@@
The Hölder’s inequality yields $\|x\|_{\ell^{q/\theta}(\ell^{p/\theta})}\leq\|x\|_{\ell^{q}(\ell^{p})}^{\theta}\|x\|_{\ell^{\infty}}^{1-\theta}$ for all $0<p,q\leq\infty$ and $0<\theta<1$. Therefore,
@@
Let $0<q_{0}<q_{1}\leq\infty$. Let $X$ be a fixed quasi-Banach space and let $X_{\bullet}=(X_{j})_{j=0}^{\infty}\subset X$ be a sequence of finite dimensional subspaces with the norm induced from $X$.
@@
(3.2) $\ell^{q_{0}}(X_{j})_{j=0}^{\infty}\hookrightarrow\ell^{q_{1}}(X_{j})_{j=0}^{\infty}.$
When $\sup_{j}\dim X_{j}=\infty$ and $q_{1}<\infty$ the map is not finitely strictly singular. For the special case $X_{j}=\ell^{p}\{1,\dots,2^{j}\}$, see Proposition 3.8 below.
@@
Let $\varepsilon>0$ and $V\subset\ell^{q_{0}}(X_{j})_{j=0}^{\infty}$ be an infinite dimensional subspace. We need to find a non-zero $x=(x_{j})_{j=0}^{\infty}\in V$ such that $\|x\|_{\ell^{q_{1}}(X)}<\varepsilon\|x\|_{\ell^{q_{0}}(X)}$.
@@
Since $q_{0}<\infty$ and $X_{j}$ are all finite dimensional, we can find out a sequence $(j_{n})_{n=1}^{N}\subset\mathbb{N}_{0}$ and a $(x^{(n)})_{n=1}^{N}\subset V$ with $\|x^{(n)}\|_{\ell^{q_{0}}(X)}=1$ for each $n\geq 1$, such that
@@
  • $\sum_{j=j_{n}+1}^{\infty}\|x^{(n)}_{j}\|_{X_{j}}^{q_{0}}\leq\delta$, or equivalently, $\sum_{j=j_{n-1}+1}^{j_{n}}\|x^{(n)}_{j}\|_{X_{j}}^{q_{0}}\geq 1-\delta$.
The construction is done by induction. We start with a unit vector $x^{(1)}\in V$, since $q_{0}<\infty$ there is a $j_{1}\geq 1$ such that $\sum_{j=j_{1}+1}^{\infty}\|x^{(1)}_{j}\|_{X_{j}}^{q_{0}}\leq\delta$.
Suppose for $n\geq 2$ and we have obtained $x^{(1)},\dots,x^{(n-1)}$ and $j_{1}<\dots<j_{n-1}$. Consider the projection $\mathcal{P}_{n-1}:\ell^{q_{0}}(\mathbb{N}_{0};X)\twoheadrightarrow\ell^{q_{0}}(\{0,\dots,j_{n-1}\};X)$, we have $\operatorname{ker}(\mathcal{P}_{n-1}|_{V})\neq\{0\}$ since $\dim\mathcal{P}_{n-1}(V)=\sum_{j=0}^{j_{n-1}}\dim X_{j}<\infty$ but $\dim V=\infty$. Therefore, there must be a unit vector $x^{(n)}\in V$ such that $\mathcal{P}_{n-1}x^{(n)}=0$, i.e. $x^{(n)}_{j}\equiv 0$ for every $0\leq j\leq j_{n-1}$. Take a $j_{n}>j_{n-1}$ be such that $\sum_{j=j_{n}+1}^{\infty}\|x^{(n)}_{j}\|_{X}^{q_{0}}\leq\delta$ we close the induction.
@@
$\|x\|_{\ell^{q_{0}}(X)}\geq\|x\|_{\ell^{q_{0}}(\{1,\dots,j_{1}\};X)}=\|x^{(1)}\|_{\ell^{q_{0}}(\{1,\dots,j_{1}\};X)}\geq(1-\delta)^{1/q_{0}}>0.$
@@
$\Big\|\sum_{k=1}^{N}x^{(k)}\Big\|_{\ell^{q_{0}}(\{j_{n-1}+1,\dots,j_{n}\};X)}^{\min(1,q_{0})}$ $=\Big\|\sum_{k=1}^{n}x^{(k)}\Big\|_{\ell^{q_{0}}(\{j_{n-1}+1,\dots,j_{n}\};X)}^{\min(1,q_{0})}$
$\geq\|x^{(n)}\Big\|_{\ell^{q_{0}}(\{j_{n-1}+1,\dots,j_{n}\};X)}^{\min(1,q_{0})}-\Big\|\sum_{k=1}^{n-1}x^{(k)}\Big\|_{\ell^{q_{0}}(\{j_{n-1}+1,\dots,j_{n}\};X)}^{\min(1,q_{0})}$
$\geq\|x^{(n)}\Big\|_{\ell^{q_{0}}(\{j_{n-1}+1,\dots,j_{n}\};X)}^{\min(1,q_{0})}-\sum_{k=1}^{n-1}\|x^{(k)}\|_{\ell^{q_{0}}(\{j_{k}+1,j_{k}+2,\dots\};X)}^{\min(1,q_{0})}$
@@
$\Big\|\sum_{k=1}^{N}x^{(k)}\Big\|_{\ell^{q_{0}}(X)}^{q_{0}}\geq\sum_{n=1}^{N}\Big\|\sum_{k=1}^{N}x^{(k)}\Big\|_{\ell^{q_{0}}(\{j_{n-1}+1,\dots,j_{n}\};X)}^{q_{0}}\geq N(1-(N+1)\delta^{\min(1,q_{0}^{-1})})^{\frac{q_{0}}{\min(1,q_{0})}}.$
On the other hand, using the property $\|x\|_{\ell^{q_{1}}(S;X)}\leq\|x\|_{\ell^{q_{0}}(S;X)}$ for every $x\in\ell^{q_{0}}(X)$ and set $S$, we have, for each $1\leq n\leq N$,
$\Big\|\sum_{k=1}^{N}x^{(k)}\Big\|_{\ell^{q_{1}}(\{j_{n-1}+1,\dots,j_{n}\};X)}^{\min(1,q_{1})}=$ $\Big\|\sum_{k=1}^{n}x^{(k)}\Big\|_{\ell^{q_{1}}(\{j_{n-1}+1,\dots,j_{n}\};X)}^{\min(1,q_{1})}$
$\leq\|x^{(n)}\Big\|_{\ell^{q_{1}}(\{j_{n-1}+1,\dots,j_{n}\};X)}^{\min(1,q_{1})}$ $+\Big\|\sum_{k=1}^{n-1}x^{(k)}\Big\|_{\ell^{q_{1}}(\{j_{n-1}+1,\dots,j_{n}\};X)}^{\min(1,q_{1})}$
$\leq\|x^{(n)}\Big\|_{\ell^{q_{0}}(\{j_{n-1}+1,\dots,j_{n}\};X)}^{\min(1,q_{1})}$ $+\sum_{k=1}^{n-1}\|x^{(k)}\|_{\ell^{q_{1}}(\{j_{k}+1,j_{k}+2,\dots\};X)}^{\min(1,q_{1})}\leq 1+N\delta^{\frac{\min(1,q_{1})}{q_{0}}}.$
@@
$\Big\|\sum_{k=1}^{N}x^{(k)}\Big\|_{\ell^{q_{1}}(\{j_{N}+1,j_{N}+2,\dots\};X)}^{\min(1,q_{1})}\leq\sum_{k=1}^{N}\|x^{(k)}\|_{\ell^{q_{1}}(\{j_{N}+1,j_{N}+2,\dots\};X)}^{\min(1,q_{1})}\leq\delta^{\frac{\min(1,q_{1})}{q_{0}}}N.$
@@
$\Big\|\sum_{k=1}^{N}x^{(k)}\Big\|_{\ell^{q_{1}}(X)}\leq\bigg(\sum_{n=1}^{N+1}\Big\|\sum_{k=1}^{N}x^{(k)}\Big\|_{\ell^{q_{1}}(\{j_{n-1}+1,\dots,j_{n}\};X)}^{q_{1}}\bigg)^{1/q_{1}}\leq N^{1/q_{1}}(1+(N+1)\delta^{\frac{\min(1,q_{1})}{q_{0}}})^{\frac{1}{\min(1,q_{1})}}.$
@@
$\frac{\|x^{(1)}+\dots+x^{(N)}\|_{\ell^{q_{1}}(X)}}{\|x^{(1)}+\dots+x^{(N)}\|_{\ell^{q_{0}}(X)}}\leq N^{\frac{1}{q_{1}}-\frac{1}{q_{0}}}\frac{(1+(N+1)\delta^{\frac{\min(1,q_{1})}{q_{0}}})^{\frac{q_{1}}{\min(1,q_{1})}}}{(1-(N+1)\delta^{\min(1,\frac{1}{q_{0}})})^{\frac{q_{0}}{\min(1,q_{0})}}}<\frac{\varepsilon}{2}\cdot 2=\varepsilon.$
@@
When $p_{0}=p_{1}$, we can take the $2^{n}$ dimensional subspace $V_{n}:=2^{-n/{p_{0}}}\cdot\ell^{p_{0}}\{1,\dots,2^{n}\}\hookrightarrow\ell^{q_{0}}(2^{-j/{p_{0}}}\cdot\ell^{p_{0}}\{1,\dots,2^{j}\})_{j=0}^{\infty}$, where $n\geq 1$. The embedding (3.4) restricted to $V_{n}$ is the identity map, hence the Bernstein numbers are all bounded from below by $1$, which does not tend to 0.
@@
Since $(S_{n},\mu_{n})$ and $(\{1,\dots,2^{n}\},2^{-n}\cdot\#)$ are both probability spaces with uniform distributions on their sample sets, we see that they are isomorphic by taking any bijection $\Gamma_{n}:\{1,\dots,2^{n}\}\to S_{n}$ between sets. Thus we have elements $r_{n,j}\circ\Gamma_{n}\in\ell^{p_{1}}\{1,\dots,2^{n}\}$.
@@
    Suppose $q_{0}<q_{1}$, and either $p_{1}\leq p_{0}<\infty$ or $p_{0}=p_{1}=\infty$, then the embedding is strictly singular but not finitely strictly singular.
@@
The map $J_{s}(x_{j})_{j=A}^{\infty}:=(2^{js}x_{j})_{j=A}^{\infty}$ defines an isomorphism $J_{s}:\ell^{q}(2^{j(s-\frac{n}{p})}\cdot\ell^{p}\{1,\dots,2^{nj}\})_{j=A}^{\infty}\to\ell^{q}(2^{-j\frac{n}{p}}\cdot\ell^{p}\{1,\dots,2^{nj}\})_{j=A}^{\infty}$ for every $0<p,q\leq\infty$. Therefore it suffices to consider the case $s=0$.
i: For $j\geq 0$ and $f:\{1,\dots,2^{nj}\}\to\mathbb{R}$ we define $\Gamma_{j}f:[0,1]\to\mathbb{R}$ by $\Gamma_{j}f(x):=f(\lfloor 2^{nj}x\rfloor)$. Clearly $\Gamma_{j}:2^{-jn/p}\cdot\ell^{p}\{1,\dots,2^{nj}\}\to L^{p}[0,1]$ are isometric embeddings for all $0<p\leq\infty$ and $j\geq 0$. Therefore $(\Gamma_{j})_{j=A}^{\infty}:\ell^{q}(2^{-jn/p}\cdot\ell^{p}\{1,\dots,2^{nj}\})_{j=A}^{\infty}\to\ell^{q}(L^{p}[0,1])$ is an isometry for all $0<p,q\leq\infty$ and $A\geq 1$.
@@
ii: The non finitely strictly singular part follows from Proposition 3.8. The strictly singular part follows from applying Proposition 3.7 to the second arrow below:
[…]
@@
Here in the lemma we have the commutative diagram $Y_{0}$$Y_{1}$$X_{0}$$X_{1}$$\scriptstyle{T_{Y}}$$\scriptstyle{T_{X}}$$\scriptstyle{\iota_{0}}$$\scriptstyle{\iota_{1}}$.
@@
Since $\iota_{0},\iota_{1}$ are both topological embeddings, there is a $C>0$ such that for $k=0,1$ and for all $x\in X_{k}$, $C^{-1}\|x\|_{X_{k}}\leq\|\iota_{k}x\|_{Y_{k}}\leq C\|x\|_{X_{k}}$. Both results on the finitely strictly singularity follow from the fact that $b_{n}(T_{X})\leq C^{2}b_{n}(T_{Y})$. The strictly singularity follows from the same argument and we leave the details to the reader. ∎
@@
$i$: When $s_{0}-s_{1}=\frac{n}{p_{0}}-\frac{n}{p_{1}}$ and $q_{0}=q_{1}$, we are going to find an infinite dimensional subspace $X$ such that the embedding restricted to $X$ has bounded inverse.
[…]
@@
For embeddings between Besov spaces with the same level of smoothness, we have:
Proposition $4.$
@@
Still we take an $\varepsilon>0$ be such that $p_{0},q_{0},p_{1},q_{1}>\varepsilon$ and $|s_{0}|,|s_{1}|<\varepsilon^{-1}$. Let $A\geq 1$ and $R_{j},S_{j}\subset Q_{j}\times G_{j}$ for $j\geq A$ be from Proposition 2.16 where we choose fixed subsets $V\Subset\Omega\Subset U$. Recall that $R_{j}=2^{(j-A)n}$ for $j\geq A$, $S_{j}=2^{(j+A)n}$ for $j\geq 0$, and for $k=0,1$ both arrows in $\ell^{q_{k}}(2^{j(s-\frac{n}{p_{k}}+\frac{n}{2})}\cdot\ell^{p_{k}}(R_{j}))_{j=A}^{\infty}\xrightarrow{\Lambda^{-1}}\mathscr{B}_{p_{k}q_{k}}^{s}(\Omega)\xrightarrow{\Lambda\circ E}\ell^{q_{k}}(2^{j(s_{k}-\frac{n}{p_{k}}+\frac{n}{2})}\cdot\ell^{p_{k}}(S_{j}))_{j=0}^{\infty}$ are bounded.
@@
iii: By Corollary 2.18 $i$ the map is noncompact. For finitely strictly singularity, when $s_{0}-s_{1}=\frac{n}{p_{0}}-\frac{n}{p_{1}}$ the result follows from Proposition 4.2 i; when $s_{0}-s_{1}=0$ the result follows from Proposition 4.4 i.
@@
v: When $s_{0}-s_{1}=\frac{n}{p_{0}}-\frac{n}{p_{1}}$ the result follows from Proposition 4.2 $i$; when $s_{0}-s_{1}=0$ the result follows from Proposition 4.4 iii. ∎